OpenStudy (anonymous):

7 sqrt 8+ 2 sqrt 12 - 3 sqrt 2

OpenStudy (anonymous):

simplify \[\sqrt{8} and \sqrt{12}\]

OpenStudy (anonymous):

\[7\sqrt{8}= 14\sqrt{2} and 2\sqrt{12}= 4\sqrt{3}\]

OpenStudy (anonymous):

what's next?

OpenStudy (anonymous):

\[(7\sqrt{8}-3\sqrt{2})+4\]

OpenStudy (anonymous):

im sooo confused. distributive property?

OpenStudy (anonymous):

associative property

OpenStudy (anonymous):

a+b+c=a+c+b

OpenStudy (anonymous):

oh... k thx

OpenStudy (anonymous):

distributive property is a(b+c)=ab+ac

OpenStudy (anonymous):

oh. i knew that ^_^